# Effects of galactic environment on size and dark matter content in low-mass galaxies

### v1

@@ 4 Summary and Conclusions
By examining these relationships across multiple data splits (500 iterations of training and validation) we find that the feature importances were stable and robust. This consistency reaffirms that environment, while not the sole driver, is a vital factor shaping galaxy sizes and inner dark matter content, particularly at lower halo masses and for galaxies more strongly perturbed by their surroundings. Additionally, our results show that other features play comparatively smaller roles, and that, at fixed stellar mass, the influence of environment can be overshadowed by mass-based factors as halo masses become large.
@@ Acknowledgments
FJM is funded by the National Science Foundation (NSF) Math and Physical Sciences (MPS) Award AST-2316748. JSB is supported by NSF grant AST-2408246 and NASA grant 80NSSC22K0827. CK is supported by NSF GRFP grant DGE-1839285 and NASA grant 80NSSC22K0827. LN is supported by the Sloan Fellowship, the NSF CAREER award 2337864, NSF award 2307788, and NSF award PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/). PFH is supported by a Simons investigator award.
We thank Joanna (Asia) Piotrowska-Karpov for her valuable input and feedback on our Random Forest analysis. We also thank Edwin J. Menendez for advising us on colorblind-friendly colormaps to use for our figures.

### v2

@@ 4 Summary and Conclusions
By examining these relationships across multiple data splits (500 iterations of training and test) we find that the feature importances were stable and robust. This consistency reaffirms that environment, while not the sole driver, is a vital factor shaping galaxy sizes and inner dark matter content, particularly at lower halo masses and for galaxies more strongly perturbed by their surroundings. Additionally, our results show that other features play comparatively smaller roles, and that, at fixed stellar mass, the influence of environment can be overshadowed by mass-based factors as halo masses become large.
@@ Acknowledgments
FJM is funded by the National Science Foundation (NSF) Math and Physical Sciences (MPS) Award AST-2316748. JSB is supported by NSF grant AST-2408246 and NASA grant 80NSSC22K0827. CK is supported by NSF Graduate Research Fellowship Program (GRFP) grant DGE-1839285 and NASA grant 80NSSC22K0827. LN is supported by the Sloan Fellowship, the NSF CAREER award 2337864, NSF award 2307788, and NSF award PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/). PFH is supported by a Simons investigator award.
We thank Edwin J. Menendez for advising us in selecting colorblind-friendly colormaps to use for our figures. Finally, we sincerely thank the referee for their thoughtful and constructive feedback, which has helped improve this manuscript.